The base of a right prism is an equilateral triangle whose edges is 10cm each and its lateral edge is 80cm.
1. Find its volume in cm3.

2. Determine the total area of the prism in cm2.
1
Expert's answer
2020-12-14T18:28:40-0500

1) V = Ah

where V is volume, A is base area and h is height

Area of Base = area of equilateral triangle = (a² * 1.732)/4

= (10² * 1.723)/4 = 43.3cm²

V = 43.3 * 80

V = 3464.10cm³

2) Atotal = 2(Abase) +3ha

Atotal = 2(43.3) + (3*80*10)

Atotal = 86.6 + 2400 = 2486.60 cm²